OpenStudy (anonymous):

Find: a) ln (e√e) b) ln (1/e^3) c) ln (e^2x) d) ln (e/e^x)

OpenStudy (anonymous):

\[\ln (e*e^{1/2})=\]

OpenStudy (anonymous):

\[\ln (e^{3/2})=3/2\ln(e)\]

OpenStudy (anonymous):

Thank you, I understood this :) :D

OpenStudy (cwrw238):

ln (e√e) = ln e + ln √e = 1 + 0.5 = 1.5

OpenStudy (anonymous):

not done ln(e)=1 so

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

what about others

OpenStudy (anonymous):

I understood it. :DThank you for your help!